If a plane can travel 460 miles per hour with the wind and 420 miles per hour agianst the wind, find the speed of the plane in still air.
Answer by stanbon(75887)

If a plane can travel 460 miles per hour with the wind and 420 miles per hour against the wind, find the speed of the plane in still air.
-------------------------------
Down-wind DATA:
p + w = 460
p - w = 420
-----------------
Add and solve for "p":
2p = 880
plane speed = 440 mph
--------
Solve for "w":
p + w = 460
440 + w = 460
w = 20 mph (wind speed)
-------------------------
Cheers,
Stan H.
=========================